During a 7.0-s time interval, a flywheel with a constant angular acceleration turns through 600 radians that acquire an angular velocity of 100 rad/s.

  1. The angular velocity at the beginning of the 7.0s is _________
  2. The angular acceleration of the flywheel is ________
1
Expert's answer
2021-03-18T20:11:36-0400

1. By definition, the angular acceleration is:


"\varepsilon =\dfrac{\Delta \omega }{\Delta t}"

where "\Delta \omega = 100 rad/s" is the gain in angular velocity in time interval "\Delta t = 7s". Thus, obtain:


"\varepsilon = \dfrac{100rad/s}{7s}\approx 14.3\space rad/s^2"

2. The number of radians the wheel turned throug is given as follows:


"\varphi = \omega_0(\Delta t) + \dfrac{\varepsilon(\Delta t)^2}{2}"

where "\omega _0" is the initial angular velocity. Substituting the expression for "\varepsilon" and "\varphi = 600rad" and expressing "\omega_0", obtain:


"\omega_0 = \dfrac{\varphi}{\Delta t}-\dfrac{\Delta \omega }{2}\\
\omega_0 = \dfrac{600rad}{7s}-\dfrac{100rad/s}{2} \approx 35.7\space rad/s"

Answer. 1) 35.7 rad/s, 2) 14.3 rad/s^2.